OpenStudy (moonlitfate):

Evaluate the expression without using a calculator. (Sketch a right triangle).

OpenStudy (moonlitfate):

\[\tan (\arccos \frac{ \sqrt{2} }{ 2 })\]

OpenStudy (anonymous):

same as before

OpenStudy (anonymous):



OpenStudy (anonymous):

find the other side you can do it in your head \[2^2-\sqrt2^2=4-2=2\]

OpenStudy (anonymous):

method 2 know tha t \[\cos(\frac{\pi}{4})=\frac{\sqrt2}{2}\] and that \[\tan(\frac{\pi}{4})=1\]

OpenStudy (moonlitfate):

APologies. I'm just trying to really understand this. Maybe, I'm thinking about it too much.

OpenStudy (anonymous):

yes you are thinking way too much i think then again if you have not seen trig for a while, who can remember all this stuff?

OpenStudy (moonlitfate):

Then again, the massive amount of Trig. and Calculus rust. Yeah, I haven't seen any of this for a good year.